# Benchmarking Sub-Genre Classification For Mainstage Dance Music

### v2

@@ 3.1.2 Annotation
Ours (ViT_B_16) $0.077$ $0.039$ $0.039$ $0.039$ $0.224$ $0.199$ $0.199$ $0.199$ $0.109$ $0.066$ $0.066$ $0.066$
Ours (VGG11_BN) $0.746$ $0.751$ $0.739$ $0.765$ $0.716$ $0.748$ $0.735$ $0.757$ $0.722$ $0.747$ $0.731$ $0.756$
Ours (DenseNet201) $0.730$ $0.767$ $0.786$ 0.790 $0.712$ $0.749$ $0.752$ $0.743$ $0.716$ $0.741$ $0.757$ $0.751$
Ours (ResNet152) $0.760$ $0.761$ $0.776$ $0.780$ $0.720$ $0.737$ $0.748$ 0.761 $0.718$ $0.744$ $0.756$ 0.764
@@ 4 Results
This section presents the experiment results alongside our interpretations and findings. Table 3 provides numerical details and Figure 2 shows the final embedding space after feature reduction by PCA, t-SNE [34] and UMAP [35]. By convention, precision, recall, and F1-score were computed after sharpening the soft labels into hard categories (e.g., converting a soft label of [$0.25,0.75,0,\ldots$] into [$0,1,0,\ldots$]) for evaluation purposes. This approach allows the results to be interpreted as the model’s ability to correctly identify the primary characteristic of a song.

### v3

@@ 3.1.2 Annotation
Ours (ViT_B_16) 0.0770.0770.077 0.0390.0390.039 0.0390.0390.039 0.0390.0390.039 0.2240.2240.224 0.1990.1990.199 0.1990.1990.199 0.1990.1990.199 0.1090.1090.109 0.0660.0660.066 0.0660.0660.066 0.0660.0660.066
Ours (VGG11_BN) 0.7460.7460.746 0.7510.7510.751 0.7390.7390.739 0.7650.7650.765 0.7160.7160.716 0.7480.7480.748 0.7350.7350.735 0.7570.7570.757 0.7220.7220.722 0.7470.7470.747 0.7310.7310.731 0.7560.7560.756
Ours (DenseNet201) 0.7300.7300.730 0.7670.7670.767 0.7860.7860.786 0.790 0.7120.7120.712 0.7490.7490.749 0.7520.7520.752 0.7430.7430.743 0.7160.7160.716 0.7410.7410.741 0.7570.7570.757 0.7510.7510.751
Ours (ResNet152) 0.7600.7600.760 0.7610.7610.761 0.7760.7760.776 0.7800.7800.780 0.7200.7200.720 0.7370.7370.737 0.7480.7480.748 0.761 0.7180.7180.718 0.7440.7440.744 0.7560.7560.756 0.764
@@ 4 Results
This section presents the experiment results alongside our interpretations and findings. Table 3 provides numerical details and Figure 2 shows the final embedding space after feature reduction by PCA, t-SNE [34] and UMAP [35]. By convention, precision, recall, and F1-score were computed after sharpening the soft labels into hard categories (e.g., converting a soft label of [0.25,0.75,0,0.25,0.75,0$\ldots$0.25 , 0.75 , 0 , …] into [$0,\ldots$ 1 , 0 , …]) for evaluation purposes. This approach allows the results to be interpreted as the model’s ability to correctly identify the primary characteristic of a song.